The base of an auditorium is in the form of an ellipse 200 ft. long and 100 ft. wide. A pin dropped near one focus can clearly be heard at the other focus. Determine the distance between the foci to the nearest tenth of a foot. Show each step of the solution.
Answer by stanbon(75887)

The base of an auditorium is in the form of an ellipse 200 ft. long and 100 ft. wide. A pin dropped near one focus can clearly be heard at the other focus. Determine the distance between the foci to the nearest tenth of a foot. Show each step of the solution.
--------------
Coordinates of foci:: (+-c,0)
a = semi-major axis = 100 ft
b = semi-minor axis = 50 ft
---
Equation::
c^2 = a^2 - b^2
c^2 = 100^2-50^2 = 7500
c = 86.60 ft
------
Ans: 2c = 2*86.6 = 173.2 ft
-----------------
Cheers,
Stan H.
---------------